40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Sep 9 20:46:28 2013, Masayuki, Update, LSC, MICH calbration 
    Reply  Tue Sep 10 17:35:50 2013, Masayuki, Manasa, Update, LSC, MICH calbration MICH_calib_loops.pngOLTF_MICHDC.pngMICH_DC_resp.pngMICH_RF_resp.pngMICH_free_noise.png
       Reply  Thu Sep 12 23:36:25 2013, Masayuki, Update, LSC, MICH calbration BSsus.PNGITMXsus.PNGITMY.PNGfree_running.PNG
          Reply  Fri Sep 13 19:22:01 2013, Masayuki, Update, LSC, MICH calbration noise.PNG
             Reply  Mon Sep 16 14:11:47 2013, rana, Update, LSC, MICH calbration Screen_Shot_2013-09-16_at_2.10.31_PM.pngScreen_Shot_2013-09-16_at_2.18.47_PM.png
                Reply  Tue Sep 17 00:50:42 2013, Masayuki, Update, LSC, MICH calbration noise.PNG
Message ID: 9127     Entry time: Thu Sep 12 23:36:25 2013     In reply to: 9121     Reply to this: 9128
Author: Masayuki 
Type: Update 
Category: LSC 
Subject: MICH calbration 

For Modelling of the OLTF, I measured the response of the BS suspension. I used the OSEM sensor for measurement. The attatchment1 is the measured TF from C1:SUS-BS_LSC_EXC to C1:SUS-BS_SUSPOS_IN1 with exciting with random force. The measured data was fitted and the resonant frequency is 1.029(±0.005) Hz and quality factor is 12.25 (± 0.2).  Additionally I did same measurement for ITMX and ITMY. The attachment 2 and 3 are the results for ITMX and ITMY. Each eigenfrequency and Q are 1.063 (±0.008) Hz and 7.33 (±0.13) (ITMX), 1.022 (±0.005) Hz and 9.41 (±0.09) (ITMY).

 After that, I locked the MICH with AS55, and measured the PSD of error signal. I compensated the that PSD by the modelled OLTF with this suspension TF and the servo TF. The result is in attachment 4. Above 1 Hz it is quite close to the previous data by Keiko (elog#6385) But below 1 Hz there is a large dip. The error signal has also this dip. I looked for a integral filter between 0.2 Hz and 1 Hz, but I connot find a such filter. And when I locked MICH with using ASDC, there was same dip at same frequency. I don't think it's true free running noise, and I will try to fix it.

I completely forgot to mention that I fitted the modelled OLTF into the measured OLTF. I used the fitted OLTF for compensation. 

 

 

Attachment 1: BSsus.PNG  35 kB  | Hide | Hide all | Show all
BSsus.PNG
Attachment 2: ITMXsus.PNG  34 kB  | Show | Hide all | Show all
Attachment 3: ITMY.PNG  35 kB  | Hide | Hide all | Show all
ITMY.PNG
Attachment 4: free_running.PNG  32 kB  | Hide | Hide all | Show all
free_running.PNG
ELOG V3.1.3-